OpenStudy (anonymous):

Find two consecutive odd numbers such that the sum of three-sevenths of the first number and one-third of the second number is equal to thirty-eight. 29 and 31 39 and 41 49 and 51

OpenStudy (nurali):

Least integer—n: 3/7x + 1/3(x + 2) = 38 9x + 7(x + 2) = 798 9x + 7x + 14 = 798 16x = 784 x = 49 Larger integer: = 49 + 2 = 51 Answer: C. 49 and 51

OpenStudy (anonymous):

nvm i know how ro do this